CRITICAL CARE PROBLEMS:

#4. During your morning assessment, you note that a patient in hypovolemic shock has a Levophed drip infusing at 7mL/h. With a concentration of Levophed 8mg/250mL, how many mcg per minute is the patient receiving?

#5. The physician orders esmolol IV drip at 50 mcg/kg/min for tachycardia. The patient weights 132 lb. The medication is supplied in 5g/500mL of D5W. Calculate the flow rate in mL/hr.

#6. The physician has ordered a nitroglycerin in drip at 10 mcg/min for a patient with chest pain. The nitroglycerin concentration is 50 mg/250mL. How many mL per hour should the nitroglycerin be infused?

#7. The physician orders procainamide IV drip at 3 mg/min for premature ventricular contractions. The medication is supplied in 2 g/500mL of NS. Calculate the flow rate in mL/hr.

#8. A patient is receiving dopamine at 6mL/hr. The concentration of dopamine is 800 mg/250mL of 0.9% NS. The patient's weight is 75 kg. Calculate the mcg/kg/min the patient is receiving.

#9. A patient is receiving dopamine at 20mL/hr. The concentration of dopamine is 800mg/250mL of 0.9% NS. The patient's weight is 70 kg. Calculate the mcg/kg/min the patient is receiving.
Answer #1

4. Flow rate 7 ml per hour or 7/60 ml per min

250 ml contains 8 mg or 8000mcg

7 ml will contain 8000/250. *7= 224 mcg

So flow rate in mcg/ min will be 224/60

3.7 mcg/min

5. 50 mcg/kg/min

for patient with 132 lbs or 60 kg, dose will be 50*60= 300mcg/min

5000mcgin 500 ml

300 mcg will be in 500/5000 *300

30 ml

so flow rate will be 30 ml per min or 180 ml/hr

6 drip rate 10 mcg/min

50 mg or 50000mcg in 250 ml

10 mcg will be in 250/50000 *10=0.05 ml

drip rate will be .05ml/min or 3 ml/hr

7. Drip rate 3mg/ min

2000mg in 500 ml

3 mg will be in 500/2000 *3= 0.75 ml

so drip rate will bs 0.75 ml/ min or 45 ml/hr